OpenStudy (goformit100):

Find the intervals in which the function fgiven by f(x) = sinx+ cosx, 0x≤2π is strictly increasing or strictly decreasing

OpenStudy (anonymous):

I have no clue how to do this lol

OpenStudy (goformit100):

Thankyou madam

OpenStudy (anonymous):

youre welcome, sorry.

OpenStudy (anonymous):

f'(x) = (sin x + cos x)' f'(x) = cos x - sin x We'll put f'(x) = 0. cos x - sin x = 0 cos x = sin x We'll divide by cos x both sides and we'll get: sin x/cos x = 1 tan x = 1 The tangent has the positive value 1 in the 1st and 3rd quadrants. x = pi/4 (1st quadrant) x = pi + pi/4 x = 5pi/4 (3rd quadrant) The extreme points of the function are: f(pi/4) = sin pi/4 + cos pi/4 = 2sqrt2/2 = sqrt2 f(5pi/4) = sin pi/4 + cos pi/4 = -2sqrt2/2 = -sqrt2 The critical points of the function are x = pi/4 and x = 5pi/4 and the extreme points are (pi/4 ; sqrt2) and (5pi/4 ; -sqrt2). @goformit100

OpenStudy (goformit100):

Thankyou @LilySwan Madam